# Learning to Translate Noise for Robust Image Denoising

### v1

@@ 5 Conclusion
We presented a novel noise translation framework for robust image denoising. Our framework allows us to effectively remove various unseen real noise, even with limited amount of training data. By employing the noise translation network, we transform arbitrary out-of-distribution (OOD) noise into Gaussian noise for which our image denoising network has learned during training. The noise translation network is designed with well-motivated loss functions and architecture, enabling effective noise translation while preserving image contents. Our experiments demonstrate that the proposed approach significantly outperforms state-of-the-art denoising models in diverse OOD real-noise benchmarks. Finally, we highlight that the generalization issue remains a critical challenge in image denoising, and our approach offers a promising solution.
Appendix:
Learning to Translate Noise for Robust Image Denoising
A Proof on Wasserstein Distance
@@
In our approach, we utilize (S.5) to formulate $\mathcal{L}_{\text{spatial}}$ in Eq. (5) and $\mathcal{L}_{\text{freq}}$ in Eq. (10), which are employed during training to explicitly transport the translated noise distribution towards the target Gaussian distribution. We refer to [30] for a detailed discussion on Wasserstein distances and optimal transport.
@@
Denoising method Metric SIDD   Poly    CC HighISO iPhone Huawei OPPO Sony Xiaomi Total Avg.
@@
Others MaskDenoising* [5] PSNR 28.66 34.56 33.87 34.61 36.54 34.89 35.30 37.89 33.46 34.42
@@
Ours Xformer-ours PSNR 38.98 38.67 37.73 40.22 42.04 39.79 40.61 44.26 36.23 39.84
SSIM 0.9549 0.9852 0.9857 0.9857 0.9816 0.9773 0.9805 0.9879 0.9748 0.9792
@@
Baseline Generalization Metric SIDD   Poly    CC HighISO iPhone Huawei OPPO Sony Xiaomi OOD Avg.
   NAFNet [6] None PSNR 39.88 37.13 35.81 38.27 40.16 37.68 39.57 43.48 34.97 38.38
SSIM 0.9594 0.9714 0.9809 0.9784 0.9700 0.9676 0.9781 0.9824 0.9681 0.9746
@@
+ Ours PSNR 39.17 38.67 37.82 39.94 41.94 39.74 40.45 44.17 36.14 39.86
SSIM 0.9566 0.9851 0.9876 0.9853 0.9805 0.9778 0.9796 0.9869 0.9745 0.9822
B Additional experimental results
B.1 Comparison with additional denoising methods
Table S.1 presents quantitative comparisons between additional image denoising methods and our approach. Generation-based methods such as DANet [34], C2N [11], and PNGAN [4] utilize the SIDD dataset for training, focusing on synthetic-to-real noise generation. MaskDenoising [5] is trained solely on Gaussian noise ($\sigma=15$), leading to notably poor performance on real-world noise datasets. CLIPDenoising [7] leverages the CLIP encoder and incorporates additional training on synthetic noise generated using Poisson-Gaussian models for sRGB denoising. However, none of these methods outperform our approach across most benchmarks, underscoring the efficacy of our noise translation framework in handling real-world noise scenarios.
Table S.3: Extended full table of Table 4 in the main paper.
@@ B.1 Comparison with additional denoising methods
$\mathcal{D}_{\theta^{*}}-test$ $\mathcal{D}_{\theta^{*}}-train$ Metric SIDD   Poly    CC HighISO iPhone Huawei OPPO Sony Xiaomi OOD Avg.
Restormer Restormer PSNR 39.08 38.74 37.60 40.06 41.62 39.68 40.55 44.12 36.14 39.81
SSIM 0.9558 0.9846 0.9861 0.9851 0.9751 0.9761 0.9794 0.9849 0.9747 0.9807
NAFNet PSNR 39.19 38.64 37.28 39.94 41.94 39.63 40.51 43.99 36.10 39.75
SSIM 0.9567 0.9851 0.9859 0.9851 0.9816 0.9767 0.9798 0.9874 0.9743 0.9820
KBNet PSNR 38.99 38.57 37.01 39.82 41.62 39.60 40.47 43.75 36.03 39.61
SSIM 0.9555 0.9841 0.9849 0.9844 0.9755 0.9760 0.9790 0.9842 0.9736 0.9809
Xformer PSNR 39.11 38.71 37.79 40.07 41.60 39.61 40.42 44.10 36.16 39.81
SSIM 0.9559 0.9851 0.9865 0.9850 0.9771 0.9759 0.9797 0.9863 0.9745 0.9813
NAFNet Restormer PSNR 39.07 38.55 37.92 39.99 41.63 39.67 40.31 43.64 36.21 39.74
SSIM 0.9558 0.9834 0.9873 0.9852 0.9755 0.9776 0.9789 0.9821 0.9748 0.9806
NAFNet PSNR 39.17 38.67 37.82 39.94 41.94 39.74 40.45 44.17 36.14 39.86
SSIM 0.9566 0.9851 0.9876 0.9853 0.9805 0.9778 0.9796 0.9869 0.9745 0.9822
KBNet PSNR 39.07 38.62 37.60 39.84 41.68 39.77 40.45 44.01 36.07 39.76
SSIM 0.9559 0.9842 0.9871 0.9848 0.9759 0.9779 0.9793 0.9842 0.9740 0.9809
Xformer PSNR 39.04 38.64 38.01 40.07 41.91 39.74 40.48 44.07 36.19 39.89
SSIM 0.9559 0.9849 0.9876 0.9854 0.9790 0.9774 0.9798 0.9855 0.9746 0.9818
KBNet Restormer PSNR 39.02 38.54 37.72 40.13 41.55 39.73 40.46 43.82 36.22 39.77
SSIM 0.9557 0.9830 0.9856 0.9851 0.9725 0.9773 0.9796 0.9818 0.9752 0.9800
NAFNet PSNR 39.15 38.63 37.71 39.94 41.91 39.71 40.49 44.20 36.12 39.84
SSIM 0.9565 0.9849 0.9865 0.9852 0.9801 0.9775 0.9798 0.9865 0.9744 0.9819
KBNet PSNR 39.06 38.57 37.59 39.83 41.63 39.71 40.46 44.04 36.04 39.73
SSIM 0.9559 0.9840 0.9859 0.9845 0.9752 0.9773 0.9794 0.9839 0.9739 0.9805
Xformer PSNR 39.05 38.64 37.76 40.11 41.76 39.70 40.51 44.05 36.21 39.84
SSIM 0.9558 0.9846 0.9859 0.9852 0.9777 0.9770 0.9802 0.9847 0.9749 0.9813
Xformer Restormer PSNR 39.04 38.55 37.71 40.19 41.98 39.67 40.48 44.05 36.27 39.86
SSIM 0.9556 0.9849 0.9852 0.9861 0.9811 0.9775 0.9800 0.9874 0.9754 0.9822
NAFNet PSNR 39.02 38.57 37.65 40.13 42.00 39.74 40.46 44.01 36.14 39.84
SSIM 0.9555 0.9850 0.9853 0.9858 0.9818 0.9776 0.9799 0.9878 0.9747 0.9822
KBNet PSNR 39.02 38.51 37.46 39.98 41.96 39.70 40.46 43.92 36.07 39.76
SSIM 0.9554 0.9847 0.9850 0.9851 0.9813 0.9775 0.9797 0.9873 0.9741 0.9818
Xformer PSNR 38.98 38.67 37.73 40.22 42.04 39.79 40.61 44.26 36.23 39.94
SSIM 0.9549 0.9852 0.9857 0.9857 0.9816 0.9773 0.9805 0.9879 0.9748 0.9824
B.2 Comparison with generalization methods
Table S.2 showcases results from our reproduction of the AFM [29] and LAN [12] methods using their officially released code, in combination with NAFNet, to enable a direct comparison with our proposed method. Applying LAN to other denoising architectures from our experiments, such as Restormer [36], KBNet [42], and Xformer [38], was infeasible due to excessive memory demands of LAN when processing images with resolutions exceeding 512 $\times$ 512. Additionally, our reproduction experiments revealed that LAN’s adaptation methodology not only requires substantial computational resources but also fails to improve generalization performance at resolutions exceeding 256 $\times$ 256. In contrast, our noise translation framework consistently delivers robust and superior generalization performance across diverse out-of-distribution (OOD) benchmarks, outperforming both AFM and LAN.
B.3 Expanded Table: Adaptability of the Noise Translation Network
Table S.3 extends the results of Table 4 in the main paper, analyzing various combinations of denoising networks used during the training and inference phases within our framework. In this context, $D_{\theta^{*}}-test$ refers to the pretrained denoising network utilized during inference, while $D_{\theta^{*}}-train$ denotes the pretrained denoising network employed for training the noise translation network. For instance, when $D_{\theta^{*}}-test$ is NAFNet and $D_{\theta^{*}}-train$ is KBNet, the noise translation network is trained to map real-world noise to align with KBNet’s prior but is evaluated with NAFNet during inference. Notably, when $D_{\theta^{*}}-test$ and $D_{\theta^{*}}-train$ correspond to different pretrained denoising models, our framework maintains high performance by effectively translating noise distributions to align with the priors of the inference model. This adaptability implies that as new denoising architectures are introduced, they can be seamlessly incorporated into our framework, potentially achieving even greater performance improvements without the need to retrain the noise translation network. Such flexibility ensures the practical applicability of our noise translation framework in real-world scenarios.
Table S.4: Effects of using Gaussian noise injection and explicit noise translation loss.
@@ B.3 Expanded Table: Adaptability of the Noise Translation Network
Metric SIDD   Poly    CC HighISO iPhone Huawei OPPO Sony Xiaomi OOD Avg.
Baseline Translation PSNR 39.35 38.32 37.25 39.22 40.80 39.24 39.75 43.86 35.74 39.27
+ Gaussian Noise Injection PSNR 39.05 38.54 37.58 39.79 41.53 39.68 40.40 43.89 36.00 39.61
+ Explicit Noise Translation Loss PSNR 39.37 37.91 37.16 39.07 40.28 37.85 39.06 42.72 35.21 38.66
SSIM 0.9574 0.9791 0.9862 0.9783 0.9650 0.9613 0.9670 0.9789 0.9605 0.9720
+ Both (Ours) PSNR 39.17 38.67 37.82 39.94 41.94 39.74 40.45 44.17 36.14 39.86
SSIM 0.9566 0.9851 0.9876 0.9853 0.9805 0.9778 0.9796 0.9869 0.9745 0.9822
@@
B.4 Expanded Table: Gaussian Injection Block and Explicit Noise Translation Loss
@@
Table S.4 provides the complete results for the OOD benchmarks in Table 3 of the main paper. These results further demonstrate the effectiveness of the Gaussian injection block and explicit noise translation loss within the proposed framework.

### v2

@@ 5 Conclusion
We presented a novel noise translation framework for robust image denoising. Our framework allows us to effectively remove various unseen real noise, even with a limited amount of training data. By employing the noise translation network, arbitrary out-of-distribution (OOD) noise is translated into Gaussian noise for which an image denoising network has learned during training. The noise translation network is designed with well-motivated loss functions and architecture, enabling effective noise translation while preserving image contents. Our experiments demonstrate that the proposed approach significantly outperforms state-of-the-art denoising models in diverse OOD real-noise benchmarks. Finally, we highlight that the generalization issue in real-world scenarios remains a critical challenge in image denoising, and our approach offers a promising solution.
Acknowledgements
This work was supported by the National Research Foundation of Korea (NRF) [RS-2022-NR070855, Trustworthy Artificial Intelligence] and the Institute of Information & Communications Technology Planning & Evaluation (IITP) [RS2022-II220959 (No.2022-0-00959), (Part 2) Few-Shot Learning of Causal Inference in Vision and Language for Decision Making; RS-2025-25442338, AI star Fellowship Support Program (Seoul National Univ.); No.RS-2021-II211343, Artificial Intelligence Graduate School Program (Seoul National University)], funded by the Korea government (MSIT).
Supplementary Material
footnotetext: indicates equal contribution.
6 Proof on Wasserstein Distance
@@
In our approach, we utilize (16) to formulate $\mathcal{L}_{\text{spatial}}$ in Eq. (5) and $\mathcal{L}_{\text{freq}}$ in Eq. (10), which are employed during training to explicitly transport the translated noise distribution towards the target Gaussian distribution. We refer to [37] for a detailed discussion on Wasserstein distances and optimal transport.
7 Additional training details
The denoising models are trained for 200K iterations with a batch size of 32, except for Restormer and Xformer, where the batch size is reduced to 4 due to the limitation of computational resources. The noise translation network is trained for 5K iterations with a batch size of 4. The explicit noise translation loss was adopted only after half (2.5k) iterations, as early translated noise exhibits large deviations that destabilize optimization. Denoising network and noise translation network are trained with the AdamW [24] optimizer with an initial learning rate of $10^{-3}$, which is reduced to $10^{-7}$ and $10^{-5}$ by a cosine annealing schedule, respectively. Each image is randomly cropped to $256\times 256$ for training. All trainings were conducted using two NVIDIA RTX A6000 GPUs.
@@ 7 Additional training details
In-distribution Out-of-distribution
Denoising Algo. Metric SIDD Poly $C$ HighISO iPhone Huawei OPPO Sony Xiaomi OOD Avg.
Restormer* [43] PSNR 40.02 37.66 36.33 38.29 40.13 38.42 39.56 44.19 35.65 38.78
SSIM 0.9603 0.9793 0.9807 0.9756 0.9734 0.9675 0.9773 0.9894 0.9710 0.9768
Restormer + NTN PSNR 39.22 38.76 37.68 40.14 41.85 39.72 40.64 44.29 36.19 39.91
SSIM 0.9569 0.9854 0.9866 0.9857 0.9786 0.9766 0.9800 0.9871 0.9750 0.9819
NAFNet* [6] PSNR 39.97 37.17 35.69 38.32 40.25 37.73 39.64 43.65 34.99 38.43
SSIM 0.9600 0.9717 0.9811 0.9788 0.9707 0.9680 0.9786 0.9829 0.9685 0.9750
NAFNet + NTN PSNR 39.24 38.72 37.84 40.00 42.08 39.83 40.55 44.34 36.17 39.94
SSIM 0.9570 0.9855 0.9877 0.9856 0.9812 0.9782 0.9801 0.9875 0.9749 0.9826
KBNet* [49] PSNR 40.35 36.82 35.23 38.09 38.00 35.18 37.80 41.79 34.25 37.15
SSIM 0.9623 0.9789 0.9810 0.9788 0.9533 0.9462 0.9663 0.9790 0.9644 0.9685
KBNet + NTN PSNR 39.13 38.62 37.61 39.89 41.76 39.80 40.57 44.22 36.07 39.82
SSIM 0.9564 0.9843 0.9861 0.9849 0.9761 0.9777 0.9798 0.9845 0.9743 0.9810
XFormer* [45] PSNR 39.98 37.45 35.95 37.86 39.99 38.35 39.61 43.89 35.54 38.58
SSIM 0.9603 0.9778 0.9782 0.9741 0.9732 0.9678 0.9779 0.9892 0.9706 0.9761
XFormer + NTN PSNR 39.10 38.75 37.82 40.29 42.20 39.89 40.70 44.44 36.25 40.04
SSIM 0.9557 0.9856 0.9862 0.9860 0.9826 0.9779 0.9807 0.9886 0.9751 0.9828
[…]
Denoising method Metric SIDD   Poly    $C$ HighISO iPhone Huawei OPPO Sony Xiaomi Total Avg.
@@
Others PDDenoising* [50] PSNR 33.95 37.10 35.82 37.09 39.13 37.69 38.75 43.60 35.03 37.57
SSIM 0.8986 0.9716 0.9742 0.9628 0.9611 0.9592 0.9707 0.9866 0.9626 0.9608
@@
MaskDenoising* [5] PSNR 28.66 34.56 33.87 34.61 36.54 34.89 35.30 37.89 33.46 34.42
IDF* [15] PSNR 31.73 37.22 36.14 37.50 40.36 37.77 38.64 42.58 34.95 37.43
SSIM 0.8011 0.9790 0.9809 0.9759 0.9776 0.9649 0.9728 0.9848 0.9617 0.9554
Ours NAFNet + NTN PSNR 39.24 38.72 37.84 40.00 42.08 39.83 40.55 44.34 36.17 39.86
SSIM 0.9570 0.9855 0.9877 0.9856 0.9812 0.9782 0.9801 0.9875 0.9749 0.9797
@@
Baseline Generalization Metric SIDD   Poly    $C$ HighISO iPhone Huawei OPPO Sony Xiaomi OOD Avg.
NAFNet [6] None PSNR 39.97 37.17 35.69 38.32 40.25 37.73 39.64 43.65 34.99 38.43
SSIM 0.9600 0.9717 0.9811 0.9788 0.9707 0.9680 0.9786 0.9829 0.9685 0.9750
@@
+ NTN PSNR 39.24 38.72 37.84 40.00 42.08 39.83 40.55 44.34 36.17 39.94
SSIM 0.9570 0.9855 0.9877 0.9856 0.9812 0.9782 0.9801 0.9875 0.9749 0.9826
8 Additional experimental results
8.1 Model-agnostic applicability
To demonstrate the model-agnostic nature of our proposed noise translation framework, we applied our method to distinct architectural paradigms, specifically the transformer-based Restormer [43] and the kernel-based KBNet [49], as noted in the main paper. Integrating our training framework into these baselines yielded consistent performance improvements in real-world denoising tasks, as detailed in Table 7. These results verify that our approach is not tailored to a specific backbone network but functions as a highly compatible framework capable of boosting the performance of diverse denoising architectures.
8.2 Comparison with additional denoising methods
Table 8 presents quantitative comparisons between additional image denoising methods and our approach. Generation-based methods such as DANet [41], C2N [12], and PNGAN [4] utilize the SIDD dataset for training, focusing on synthetic-to-real noise generation. PDDenoising [50] applies pixel-shuffle down-sampling to denoise complex real noise with a simple Gaussian denoiser. However, pixel-shuffling alters image semantics, and optimal shuffling parameters vary per image, making optimization challenging and yielding poor performance. MaskDenoising [5] and IDF [15] are trained solely on Gaussian noise ($\sigma=15$), leading to notably poor performance on real-world noise datasets. CLIPDenoising [8] leverages the CLIP encoder and incorporates additional training on synthetic noise generated using Poisson-Gaussian models for sRGB denoising. However, none of these methods outperform our approach across most benchmarks, underscoring the efficacy of our noise translation framework in handling real-world noise scenarios.
Table 10: Expanded table of Table 2: Gaussian injection and explicit noise translation loss.
@@ 8.2 Comparison with additional denoising methods
Metric SIDD Poly $C$ HighISO iPhone Huawei OPPO Sony Xiaomi OOD Avg.
w/o GIBlock PSNR 39.35 38.32 37.25 39.22 40.80 39.24 39.75 43.86 35.74 39.27
w/o Explicit Loss PSNR 39.05 38.54 37.58 39.79 41.53 39.68 40.40 43.89 36.00 39.61
w/o adaptive gating ($\mathcal{L}_{\text{explicit}}$) PSNR 39.33 38.59 37.10 39.35 41.69 39.56 40.27 44.34 35.95 39.61
SSIM 0.9572 0.9846 0.9864 0.9821 0.9793 0.9753 0.9782 0.9888 0.9720 0.9808
Full (Ours) PSNR 39.24 38.72 37.84 40.00 42.08 39.83 40.55 44.34 36.17 39.94
SSIM 0.9570 0.9855 0.9877 0.9856 0.9812 0.9782 0.9801 0.9875 0.9749 0.9826
Table 11: Expanded table of Table 4: Adaptability of the noise translation network.
@@
$\mathcal{D}_{\theta^{*}}$ for testing $\mathcal{D}_{\theta^{*}}$ for training $\mathcal{T}_{\phi}$ Metric SIDD   Poly    $C$ HighISO iPhone Huawei OPPO Sony Xiaomi OOD Avg.
Restormer Restormer PSNR 39.22 38.76 37.68 40.14 41.85 39.72 40.64 44.29 36.19 39.91
SSIM 0.9569 0.9854 0.9866 0.9857 0.9786 0.9766 0.9800 0.9871 0.9750 0.9819
NAFNet PSNR 39.16 38.65 37.34 39.98 41.92 39.73 40.56 43.95 36.19 39.79
SSIM 0.9564 0.9851 0.9861 0.9854 0.9811 0.9770 0.9800 0.9871 0.9750 0.9821
KBNet PSNR 39.21 38.53 37.05 39.88 41.93 39.65 40.53 43.81 36.14 39.69
SSIM 0.9567 0.9848 0.9851 0.9850 0.9817 0.9766 0.9799 0.9873 0.9743 0.9819
Xformer PSNR 39.12 38.73 37.80 40.11 41.65 39.64 40.42 44.11 36.18 39.83
SSIM 0.9559 0.9850 0.9867 0.9852 0.9774 0.9763 0.9795 0.9866 0.9746 0.9814
NAFNet Restormer PSNR 39.16 38.61 38.00 40.07 41.95 39.71 40.41 44.05 36.22 39.88
SSIM 0.9565 0.9846 0.9877 0.9856 0.9800 0.9775 0.9793 0.9858 0.9748 0.9819
NAFNet PSNR 39.24 38.72 37.84 40.00 42.08 39.83 40.55 44.34 36.17 39.94
SSIM 0.9570 0.9855 0.9877 0.9856 0.9812 0.9782 0.9801 0.9875 0.9749 0.9826
KBNet PSNR 39.17 38.67 37.54 39.83 41.95 39.73 40.45 44.18 36.11 39.81
SSIM 0.9565 0.9853 0.9870 0.9849 0.9814 0.9777 0.9798 0.9875 0.9741 0.9822
Xformer PSNR 39.06 38.66 38.04 40.07 41.93 39.72 40.46 44.11 36.17 39.89
SSIM 0.9560 0.9848 0.9877 0.9852 0.9792 0.9773 0.9794 0.9857 0.9743 0.9817
KBNet Restormer PSNR 39.14 38.59 37.79 40.14 41.77 39.67 40.45 44.02 36.22 39.83
SSIM 0.9564 0.9843 0.9863 0.9855 0.9776 0.9773 0.9798 0.9850 0.9751 0.9814
NAFNet PSNR 39.12 38.62 37.67 39.95 41.88 39.73 40.51 44.16 36.16 39.84
SSIM 0.9563 0.9849 0.9863 0.9853 0.9797 0.9776 0.9800 0.9861 0.9747 0.9818
KBNet PSNR 39.13 38.62 37.61 39.89 41.76 39.80 40.57 44.22 36.07 39.82
SSIM 0.9564 0.9843 0.9861 0.9849 0.9761 0.9777 0.9798 0.9845 0.9743 0.9810
Xformer PSNR 39.08 38.64 37.74 40.08 41.76 39.69 40.48 44.09 36.18 39.83
SSIM 0.9560 0.9845 0.9862 0.9851 0.9776 0.9772 0.9797 0.9849 0.9747 0.9812
Xformer Restormer PSNR 39.09 38.60 37.81 40.25 42.12 39.72 40.51 44.10 36.28 39.92
SSIM 0.9558 0.9851 0.9859 0.9861 0.9831 0.9776 0.9801 0.9885 0.9753 0.9827
NAFNet PSNR 39.17 38.62 37.73 40.18 42.29 39.75 40.56 44.27 36.19 39.95
SSIM 0.9563 0.9856 0.9860 0.9860 0.9851 0.9775 0.9806 0.9896 0.9750 0.9832
KBNet PSNR 39.10 38.51 37.43 39.98 42.12 39.69 40.47 43.98 36.15 39.79
SSIM 0.9557 0.9851 0.9854 0.9852 0.9844 0.9774 0.9801 0.9889 0.9743 0.9826
Xformer PSNR 39.10 38.75 37.82 40.29 42.20 39.89 40.70 44.44 36.25 40.04
SSIM 0.9557 0.9856 0.9862 0.9860 0.9826 0.9779 0.9807 0.9886 0.9751 0.9828
8.3 Comparison with generalization methods
Table 9 showcases results from our reproduction of the AFM [34] and LAN [14] methods using their officially released code, in combination with NAFNet, to enable a direct comparison with our proposed method. Applying LAN to other denoising architectures from our experiments, such as Restormer [43], KBNet [49], and Xformer [45], was infeasible due to excessive memory demands of LAN when processing images with resolutions exceeding 512 $\times$ 512. Additionally, our reproduction experiments revealed that LAN’s adaptation methodology not only requires substantial computational resources but also fails to improve generalization performance at resolutions exceeding 256 $\times$ 256. In contrast, our noise translation framework consistently delivers robust and superior generalization performance across diverse out-of-distribution benchmarks, outperforming both AFM and LAN.
@@ 8.3 Comparison with generalization methods
8.4 Expanded Table: Gaussian Injection Block and Explicit Noise Translation Loss
@@
Table 10 provides the complete results for the OOD benchmarks in Table 2 of the main paper. These results further demonstrate the effectiveness of the Gaussian injection block and explicit noise translation loss within the proposed framework.
8.5 Expanded Table: Adaptability of the Noise Translation Network
Table 11 extends the results of Table 4 in the main paper, analyzing various combinations of denoising networks used during the training and inference phases within our framework. In this context, $\mathcal{D}_{\theta^{*}}$ for testing refers to the pretrained denoising network utilized during inference, while $\mathcal{D}_{\theta^{*}}$ for training $\mathcal{T}_{\phi}$ denotes the pretrained denoising network employed for training the noise translation network. For instance, when $\mathcal{D}_{\theta^{*}}$ for testing is NAFNet and $\mathcal{D}_{\theta^{*}}$ for training $\mathcal{T}_{\phi}$ is KBNet, the noise translation network is trained to map real-world noise to align with KBNet’s prior but is evaluated with NAFNet during inference. Notably, when $D_{\theta^{*}}$ for testing and $\mathcal{D}_{\theta^{*}}$ for training $\mathcal{T}_{\phi}$ correspond to different pretrained denoising models, our framework maintains high performance by effectively translating noise distributions to align with the priors of the inference model. This adaptability implies that as new denoising architectures are introduced, they can be seamlessly incorporated into our framework, potentially achieving even greater performance improvements without the need to retrain the noise translation network. Such flexibility ensures the practical applicability of our noise translation framework in real-world scenarios.